OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = 4. P(x) = x2 - 4x + 5 P(4) = _____ Enter a number only.

OpenStudy (paki):

just put "4" in the equation and solve... @fani1996

OpenStudy (anonymous):

and tht will give my answer?

OpenStudy (paki):

yeah

OpenStudy (anonymous):

-3

OpenStudy (paki):

how "-3"... can you please show your work...?

OpenStudy (anonymous):

well i did what u told me and got tht

OpenStudy (paki):

agree... but how you got "-3"....?

OpenStudy (anonymous):

-8-16+5

OpenStudy (anonymous):

8

OpenStudy (paki):

P(x) = x2 - 4x + 5 put p(4)... P(4) = (4)^2 - 4(4) + 5.......... Now solve....

OpenStudy (anonymous):

5

OpenStudy (paki):

yeah... rite....

OpenStudy (anonymous):

:) thanks for explaining to me i forgot i was squaring aja :)

OpenStudy (paki):

pleasure...